# Optimizing Noise for $f$-Differential Privacy via Anti-Concentration and Stochastic Dominance

### v2

@@ 2 Differential Privacy Background
Let $f$ be a symmetric tradeoff function. Then
@@
Invariance to Postprocessing If $M:\mathscr{D}\rightarrow\mathscr{Y}$ satisfies $f$-DP and $T:\mathscr{Y}\rightarrow\mathscr{Z}$ is a (possibly randomized) function, then $T\circ M:\mathscr{D}\rightarrow Z$ also satisfies $f$-DP. This property is important to ensure that the privacy guarantee cannot be “undone” by a data-independent attack.
@@ 4.3 CNDs are Sub-Exponential
In the right plot of Figure 3 we compare $P(|N_{i}|\leq t)$ where $N_{2}\sim\mathrm{Cauchy}(0,1)$ and $N_{1}$ is drawn from the CND for $C_{1}$ constructed in Proposition 9. Empirically, for $t\geq 1/2$, $P(|N_{1}|\geq t)\geq P(|N_{2}|\geq t)$ and the difference is at most .$\approx.00425$ for $t\leq 1/2$. Thus, the CND would be preferred over the Cauchy distribution in most applications. To illustrate the difference in the tails, we simulated 100 random variables from both distributions [^2] The code for this paper, available at https://github.com/JordanAwan/OptimizingNoiseForFDP, gives an implementation to simulate from the CND constructed in Proposition 9.. The maximum absolute value of the CND variables was 4.28 whereas the maximum absolute value of the Cauchy random variables was 118.43. We see that the subexponential tails of the CND significantly reduce the chances of observing extreme events.
@@ 6 Discussion
The investigations of this paper were limited to univariate noise distributions. Future researchers may develop a multivariate extension to Lemma 3, and explore whether there are optimality properties of the multivariate CNDs proposed by Awan and Dong (2022).
@@
We are thankful to Zhanyu Wang for feedback and discussion on an early draft of this manuscript. We are also grateful to the anonymous reviewers, whose feedback has greatly improved the presentation of this paper. This work was supported in part by the National Science Foundation, Grant No. SES-2150615 to Purdue University.
@@ Appendix A Proofs and Technical Details
Case 1: If $f_{t}(1/2)=0$, then by symmetry of $f_{t}$, we have that $f_{t}(1)\leq 1/2$. This is because if $f=T(P,Q)=T(Q,P)$, then the type I $\alpha$ and type II error $\beta$ can be interchanged: $\beta=f(1-\alpha)$ if and only if $\alpha=f(1-\beta)$. It follows that

### v3

@@ 2 Differential Privacy Background
Let $f$ be a symmetric tradeoff function. Then,
@@
Invariance to Postprocessing: If $M:\mathscr{D}\rightarrow\mathscr{Y}$ satisfies $f$-DP and $T:\mathscr{Y}\rightarrow\mathscr{Z}$ is a (possibly randomized) function, then $T\circ M:\mathscr{D}\rightarrow Z$ also satisfies $f$-DP. This property is important to ensure that the privacy guarantee cannot be “undone” by a data-independent attack.
@@ 4.3 CNDs are Sub-Exponential
In the right plot of Figure 3 we compare $P(|N_{i}|\leq t)$ where $N_{2}\sim\mathrm{Cauchy}(0,1)$ and $N_{1}$ is drawn from the CND for $C_{1}$ constructed in Proposition 9. Empirically, for $t\geq 1/2$, $P(|N_{1}|\geq t)\geq P(|N_{2}|\geq t)$ and the difference is at most .$\approx.00425$ for $t\leq 1/2$. Thus, the CND would be preferred over the Cauchy distribution in most applications. To illustrate the difference in the tails, we simulated 100 random variables from both distributions222 The code for this paper, available at https://github.com/JordanAwan/OptimizingNoiseForFDP, gives an implementation to simulate from the CND constructed in Proposition 9.. The maximum absolute value of the CND variables was 4.28 whereas the maximum absolute value of the Cauchy random variables was 118.43. We see that the subexponential tails of the CND significantly reduce the chances of observing extreme events.
@@ 6 Discussion
Multivariate Extension: The results of this paper are limited to univariate noise distributions. Future researchers may develop a multivariate extension to Lemma 3, and explore whether there are optimality properties of the multivariate CNDs proposed by Awan and Dong (2022).
[…]
$P(\lVert N\rVert\leq t/2)$ $\leq\inf_{\lVert v\rVert=1}P(|\mathrm{Proj}_{v}N|\leq t/2)$ (1)
$\leq\inf_{\lVert v\rVert=1}\mathrm{TV}(\mathrm{Proj}_{v}N,\mathrm{Proj}_{v}N+t)$ (2)
$\leq\inf_{\lVert v\rVert=1}\mathrm{TV}(N,N+tv)$ (3)
$\leq\begin{cases}1-2f^{\circ k}(c_{f})&\text{if $t=2k+1$}\\ 1-2f^{\circ k}(1/2)&\text{if $t=2k$},\end{cases}$ (4)
where $\mathrm{Proj}_{v}$ is the orthogonal projection onto the subspace spanned by the vector $v$, (2) applies Lemma 3, (3) applies the data processing inequality for total variation, and (4) uses the same argument as in the proof of Theorem 5. While this result does give some indication of the concentration of the distribution of $N$, it is generally not tight as the bound does not involve the dimension. The following example illustrates this gap in the case of Gaussian noise.
Example 13
The distribution $N(0,I)$ is a multivariate CND with respect to $\lVert\cdot\rVert_{\infty}$ for $G_{\sqrt{d}}$-DP (Awan and Dong, 2022). We can directly calculate $P(\lVert N\rVert_{\infty}\leq t/2)=[1-2\Phi(-t/2)]^{d}$. On the other hand, the bound (3) simplifies to $P(\lVert N\rVert_{\infty}\leq t/2)\leq 1-2\Phi(-t/2),$ and we see that the bound is tight only in the case that $d=1$.
We believe that the reason our method does not produce a tight bound in multivariate settings is because it is based on a two-point hypothesis test, similar in spirit to the Le Cam method for minimax lower bounds. It is possible that using a packing argument, similar to the Fano or Assouad methods, could produce a tighter bound that incorporates the dimension.
@@
We are thankful to Zhanyu Wang for feedback and discussion on an early draft of this manuscript. This work was supported in part by the National Science Foundation, Grant No. SES-2150615 to Purdue University.
@@ Appendix A Proofs and Technical Details
Case 1: If $f_{t}(1/2)=0$, then by symmetry of $f_{t}$, we have that $f_{t}(1)\leq 1/2$. This is because if $f=T(P,Q)=T(Q,P)$, then the type I error $\alpha$ and type II error $\beta$ can be interchanged: $\beta=f(1-\alpha)$ if and only if $\alpha=f(1-\beta)$. It follows that